# The Landscape of Arabic Large Language Models (ALLMs): A New Era for Arabic Language Technology

### v1

@@
Shahad Al-Khalifa shahadalkhalifa90@gmail.com is a Researcher at iWAN Research Group, King Saud University Riyadh Riyadh Saudi Arabia Nadir Durrani is a Senior Scientist at Qatar Computing Research Institute Doha Doha Qatar ndurrani@hbku.edu.qa Hend Al-Khalifa is a Professor at King Saud University and Head of iWAN Research Group Riyadh Saudi Arabia hendk@ksu.edu.sa  and  Firoj Alam is a Senior Scientist at Qatar Computing Research Institute Doha Doha Qatar fialam@hbku.edu.qa
@@ 1. Foundations of Arabic NLP
The story of Arabic NLP began in 1985 when pioneers like Sakhr Software [^1] http://www.sakhr.com/index.php/en/ tackled the unique challenges posed by Arabic’s rich morphology and complex syntax. Early systems, such as morphological analyzers, laid the groundwork for computational tools by addressing tasks like word segmentation and root extraction—critical for processing a language with intricate grammatical structures. As illustrated in Figure 1, the early 2000s saw the rise of statistical models, with techniques like n-grams and word stemming being widely utilized for various NLP tasks such as text classification, information retrieval, and machine translation. These models offered improvements over rule-based approaches but were constrained by limited data availability and struggled to generalize across Arabic’s diverse dialects and linguistic complexities. Despite these challenges, statistical methods provided a stepping stone for future innovations, setting the stage for more advanced approaches. The 2010s marked a paradigm shift with the adoption of deep language models, bringing with them powerful tools like word embeddings and Farasa Arabic word processing tool (Abdelali et al., 2016), [^2] https://farasa.qcri.org/ which significantly enhanced the accuracy and adaptability of NLP systems. Techniques like LSTMs and CNNs enabled breakthroughs in sentiment analysis, machine translation, and dialect identification, allowing for more nuanced understanding of Arabic text. However, the inherent diversity of Arabic, including its numerous dialects and morphological richness, continued to pose challenges, underscoring the need for even more sophisticated and scalable models.
@@ 2. The Rise of Transformers and ALLMs
Among the most influential transformer-based models was BERT (Bidirectional Encoder Representations from Transformers), which revolutionized NLP by setting new benchmarks in understanding language nuances. Building on this success, specialized Arabic models like AraBERT (Antoun et al., 2020) and QARIB (Abdelali et al., 2021) were developed, significantly improving performance in tasks such as sentiment analysis, named entity recognition, and dialect identification. Additionally, tools such as CAMeL [^3] https://github.com/CAMeL-Lab/camel_tools and Farasa [^4] https://farasa.qcri.org/ offered support for various Arabic language processing tasks. These models became essential tools across a wide range of applications, showcasing the transformative potential of BERT-inspired architectures in Arabic NLP.
@@ 3. Datasets and Benchmarks for ALLMs
SFT. Instruction tuning is essential for enabling an LLM to engage in dialogue-style interactions with users. Across all ALLM initiatives, the curation of SFT datasets often began with publicly available English datasets (e.g., Super-NaturalInstructions, [^6] https://huggingface.co/allenai/open-instruct-sni-13b Natural Questions, [^7] https://huggingface.co/datasets/google-research-datasets/natural_questions P3, [^8] https://huggingface.co/datasets/bigscience/P3 xP3999https://huggingface.co/datasets/bigscience/xP3), which were subsequently translated into Arabic (Sengupta et al., 2023; Team et al., 2025). Some initiatives also developed their own in-house datasets. For instance, Jais created NativeQA, a set of question–answer pairs focused on the UAE and the surrounding region, as well as SafetyQA and DoNotAnswer, to ensure the model avoids engaging in unsafe conversations, including discussions on self-harm, sexual violence, or identity attacks. For cultural alignment, relevant efforts include the development of CIDER to ensure alignment with Arabic norms (Alyafeai et al., 2024).

### v2

@@
Shahad Al-Khalifa shahadalkhalifa90@gmail.com is a Researcher at iWAN Research Group, King Saud University Riyadh Riyadh Saudi Arabia , Nadir Durrani is a Senior Scientist at Qatar Computing Research Institute Doha Doha Qatar ndurrani@hbku.edu.qa , Hend Al-Khalifa is a Professor at King Saud University and Head of iWAN Research Group Riyadh Saudi Arabia hendk@ksu.edu.sa and Firoj Alam is a Senior Scientist at Qatar Computing Research Institute Doha Doha Qatar fialam@hbku.edu.qa
@@ 1. Foundations of Arabic NLP
The story of Arabic NLP began in 1985 when pioneers like Sakhr Software [^1] http://www.sakhr.com/index.php/en/ tackled the unique challenges posed by Arabic’s rich morphology and complex syntax. Early systems, such as morphological analyzers, laid the groundwork for computational tools by addressing tasks like word segmentation and root extraction—critical for processing a language with intricate grammatical structures. As illustrated in Figure 1, the early 2000s saw the rise of statistical models, with techniques like n-grams and word stemming being widely utilized for various NLP tasks such as text classification, information retrieval, and machine translation. These models offered improvements over rule-based approaches but were constrained by limited data availability and struggled to generalize across Arabic’s diverse dialects and linguistic complexities. Despite these challenges, statistical methods provided a stepping stone for future innovations, setting the stage for more advanced approaches. The 2010s marked a paradigm shift with the adoption of deep language models, bringing with them powerful tools like word embeddings (Soliman et al., 2017) and Arabic word processing tools like MADAMIRA (Pasha et al., 2014) and FARASA (Abdelali et al., 2016), which significantly enhanced the accuracy and adaptability of NLP systems. Techniques like LSTMs and CNNs enabled breakthroughs in sentiment analysis, machine translation, and dialect identification, allowing for more nuanced understanding of Arabic text. However, the inherent diversity of Arabic, including its numerous dialects and morphological richness, continued to pose challenges, underscoring the need for even more sophisticated and scalable models.
@@ 2. The Rise of Transformers and ALLMs
Among the most influential transformer-based models was BERT (Bidirectional Encoder Representations from Transformers), which revolutionized NLP by setting new benchmarks in understanding language nuances. Building on this success, specialized Arabic models like AraBERT (Antoun et al., 2020), CAMEL-BERT  (Inoue et al., 2021) and QARIB (Abdelali et al., 2021) were developed, significantly improving performance in tasks such as sentiment analysis, named entity recognition, and dialect identification. Additionally, tools such as CAMeL (Obeid et al., 2020) and Farasa (Abdelali et al., 2016) offered support for various Arabic language processing tasks. These models became essential tools across a wide range of applications, showcasing the transformative potential of BERT-inspired architectures in Arabic NLP.
@@ 3. Datasets and Benchmarks for ALLMs
SFT. Instruction tuning is essential for enabling an LLM to engage in dialogue-style interactions with users. Across all ALLM initiatives, the curation of SFT datasets often began with publicly available English datasets (e.g., Super-NaturalInstructions, [^3] https://huggingface.co/allenai/open-instruct-sni-13b Natural Questions, [^4] https://huggingface.co/datasets/google-research-datasets/natural_questions P3, [^5] https://huggingface.co/datasets/bigscience/P3 xP3666https://huggingface.co/datasets/bigscience/xP3), which were subsequently translated into Arabic (Sengupta et al., 2023; Team et al., 2025). Some initiatives also developed their own in-house datasets. For instance, Jais created NativeQA, a set of question–answer pairs focused on the UAE and the surrounding region, as well as SafetyQA and DoNotAnswer, to ensure the model avoids engaging in unsafe conversations, including discussions on self-harm, sexual violence, or identity attacks. For cultural alignment, relevant efforts include the development of CIDER to ensure alignment with Arabic norms (Alyafeai et al., 2024).